OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4.

OpenStudy (anonymous):

parallel lines has equal slopes thus you have y = -3x - 4 here the slope = m1 = -3 the slope of line parallel to this line is also equals m2 = -3 now you have slope = -3 and point = (2, 6) use point slope form to get the equation point slope form is y-y1 = m(x-x1) here x1 = 2 and y1 = 6 make the equation now

OpenStudy (anonymous):

Uhm y = -3x + 12

OpenStudy (anonymous):

yes exactly you are right

OpenStudy (anonymous):

Choose the equation of the vertical line passing through the point (-4, 2). y = -4 x = -4 y = 2 x = 2

OpenStudy (anonymous):

for a vertical line the value of x never changes. can you guess the right option now???

OpenStudy (anonymous):

x = -4 is the option.

OpenStudy (anonymous):



OpenStudy (anonymous):

Choose the equation of the horizontal line that passes through the point (-8, -7)